OpenStudy (anonymous):

use factoring to solve this equation x^2=25/4

OpenStudy (anonymous):

x=5/2

OpenStudy (angela210793):

x=sqr(25/4)=+-5/2

OpenStudy (anonymous):

@ishaan94. You seem to neglect the negative root.